# Controlling superradiant phase transition in quantum Rabi model

### v1

@@ IV Excitation spectra
To figure out the ground-state property of the system in the whole coupling-parameter space, we calculate the excitation energy in the limit $\tilde{\eta}\rightarrow\infty$ in this section.
@@ IV.1 Normal phase
Then, in the limit $\tilde{\eta}\rightarrow\infty$, we make an unitary transformation $U_{\text{N}}^{\dagger}{H}_{\text{AR}}U_{\text{N}}$ on Hamiltonian (16) with
@@
The transformed Hamiltonian becomes free of coupling between qubit subspaces $\mathcal{H}_{g}$ and $\mathcal{H}_{e}$, which are spanned by $\left|g\right\rangle$ and $\left|e\right\rangle$, respectively. After a projection into $\mathcal{H}_{g}$, i.e., $\tilde{H}_{\text{N}}=\langle g|U_{\text{N}}^{\dagger}H_{\text{AR}}U_{\text{N}}|g\rangle$, we obtain an effective low-energy Hamiltonian as
@@
where a constant term is neglected. Here,
Figure 3: (Color online) The phase transition boundary $\tilde{g}_{C}/\omega_{0}$ for the anisotropic QRM versus $\xi$ for $\nu/\omega_{0}=0.402$ and $0.68$ with $\eta=10^{2}$.
$\zeta=\frac{\lambda(1+\epsilon)}{2}\ \text{and}\ \zeta^{\prime}=\frac{\lambda(1-\epsilon)}{2}$ (20)
with the dimensionless rotating-coupling strength $\lambda\equiv{g_{r}/g_{C}}$. We note that there is no interaction term in Eq. (19). The excitation energy $\omega_{\text{N}}$ can be directly obtained as
@@
$\tilde{H}_{\text{N}}$ implies divergent values of $\langle x\rangle$ ($\langle p\rangle$) due to the disappearance of $x^{2}$ ($p^{2}$) term when $\epsilon>0$ ($\epsilon<0$) in the limit $\tilde{\eta}\rightarrow\infty$. By this fact, the normal phase is restricted in the region
@@
In the normal phase, the value of $\langle x\rangle\ (\langle p\rangle)$ is zero. $\tilde{g}_{C}$ determines the phase transition boundary for the anisotropic QRM. Note that, $g_{C}$ is the critical coupling strength for the standard QRM. When $g_{r}<\left|\tilde{g}_{C}\right|$, the ground state is in the normal phase with the cavity field in vacuum and qubit in $\left|g\right\rangle$. While outside the normal phase, namely, $g_{r}>\left|\tilde{g}_{C}\right|$, the system is in the superradiant phase with both the qubit and the cavity field excited.
In Fig. 3, we plot $\tilde{g}_{C}/\omega_{0}$ versus $\xi$ for $\nu/\omega_{0}=0.402\ \text{and}\ 0.68$. For $\nu/\omega_{0}=0.402$, $\tilde{g}_{C}/\omega_{0}$ oscillates between 0 and 0.0309031. The zero points of $\tilde{g}_{C}/\omega_{0}$ locate at those zeros of $J_{n_{0}}(\xi)$. The first zero point of $\tilde{g}_{C}/\omega_{0}$ locates at $\xi=5.13562$. For a larger $\nu$, such as $\nu/\omega_{0}=0.68$, $\tilde{g}_{C}/\omega_{0}$ becomes independent of $\xi$ due to $m_{0}=n_{0}$. Thus, we can tune the QPT boundary via $\nu$ and $\xi$.
@@ IV.2 Superradiant phase
Now, we study the excitation energy in the superradiant phase, namely, $g_{r}>\left|\tilde{g}_{C}\right|$. In the superradiant phase, there are four different superradiant phases. When $\epsilon>0$, the superradiant phase can be defined as superradiant X (SX) phase since the $\langle x\rangle$ becomes divergent. When $\epsilon<0$, the system is in superradiant P (SP) phase since the $\langle p\rangle$ is divergent. For $g_{cr}>2\left|g_{C}\right|$ with $g_{r}=0$, the ground state of system is in superradiant XP a (SXPa) phase while for $g_{r}>2\left|g_{C}\right|$ with $g_{cr}=0$, the system is in superradiant XP b (SXPb) phase. When $g_{cr}=0$, only rotating terms exist in the Hamiltonian, which reduces to a Jaynes-Cummings model. In the following, we investigate the excitation energy in the limit $\tilde{\eta}\rightarrow\infty$ for SX, SP, SXPa and SXPb phase, respectively. To effectively describe the superradiant phase with macroscopic excitation, we displace the cavity field $a$ by displacing operator $D_{k}(\alpha)$ (k=SX, SP, SXPa, SXPb) for k phase. After the displacement, the Hamiltonian $H_{\text{AR}}$ becomes
@@
Here, the detailed form of $D_{k}(\alpha)$ is given in the Appendix A. To diagonalize the qubit part in Hamiltonian $\tilde{H}_{AR}$, we introduce a new set of Pauli operators $\tau_{x,y,z}^{k}$ (k=SX, SP, SXPa, SXPb) for each k phase, respectively. To be concise, the form of $\tau_{x,y,z}^{k}$ for k phase is also given in the Appendix A. In terms of the new Pauli operators $\tau_{x,y,z}^{k}$ , the displaced Hamiltonian $\tilde{H}_{\text{AR}}$ becomes
@@
$\tilde{H}_{\text{AR}}(\alpha_{\text{SXPa}})=$ $\frac{\Omega_{\text{SXPa}}}{2}\tau_{z}^{{}_{\text{SXPa}}}+\chi_{3}(\frac{4}{\mu}\tau_{x}^{{}_{\text{SXPa}}}x+\mu\tau_{y}^{{}_{\text{SXPa}}}p)$
@@
$\tilde{H}_{\text{AR}}(\alpha_{\text{SXPb}})=$ $\frac{\Omega_{\text{SXPb}}}{2}\tau_{z}^{{}_{\text{SXPb}}}+\chi_{3}(\frac{4}{\lambda}\tau_{x}^{{}_{\text{SXPb}}}x-\lambda\tau_{y}^{{}_{\text{SXPb}}}p)$
@@
and the rescaled qubit frequency $\Omega_{k}$ for k phase is
@@
The displacement parameter $\alpha$ is determined by the vanishment of the non-diagonal terms during deriving Eq. (25), which is obtained as
$\alpha_{\text{SX}}$ =±$=\pm\sqrt{\frac{\tilde{\eta}}{2}(\zeta^{2}-\frac{1}{\zeta^{2}})},$ (28)
$\alpha_{\text{SP}}$ =±$=\pm\sqrt{\frac{\tilde{\eta}}{2}(\zeta^{\prime 2}-\frac{1}{\zeta^{\prime 2}})},$
$\alpha_{\text{SXPa}}$ =±$=\pm\sqrt{\tilde{\eta}(\frac{\mu^{2}}{16}-\frac{1}{\mu^{2}})},$
$\alpha_{\text{SXPb}}$ =±$=\pm\sqrt{\tilde{\eta}(\frac{\lambda^{2}}{16}-\frac{1}{\lambda^{2}})},$
for k (k=SX, SP, SXPa, SXPb) phase. To distinguish $\alpha$ in each k phase, here we have added a subscript “k” in $\alpha$, i.e., we label $\alpha$ as $\alpha_{k}$. Here, the dimensionless CR coupling strength $\mu$ is defined as $\mu=g_{cr}/g_{C}$. The parameter $\chi_{i}$ (i=1,2,3) in Eq. (25) is defined as
@@
Thus, we can make similar unitary transformation $U_{k}$ as $U_{\text{N}}$ in the normal phase on $\tilde{H}_{\text{AR}}(\alpha_{k})$ to eliminate the non-diagonal terms, namely,
@@
(k=SX, SP, SXPa, SXPb). Here, the detailed form of $U_{k}$ is given in the Appendix A. After performing the similar procedure used to derive $\tilde{H}_{\text{N}}$, we obtain the effective Hamiltonian $\tilde{H}_{k}$ in k phase as following:
$\tilde{H}_{\text{SX}}$ $=\frac{\tilde{\omega}_{c}}{2}(1-\frac{1}{\zeta^{4}})x^{2}+\frac{\tilde{\omega}_{c}}{2}(1-\frac{\zeta^{\prime 2}}{\zeta^{2}})p^{2},$ (31)
$\tilde{H}_{\text{SP}}$ $=\frac{\tilde{\omega}_{c}}{2}(1-\frac{\zeta^{2}}{\zeta^{\prime 2}})x^{2}+\frac{\tilde{\omega}_{c}}{2}(1-\frac{1}{\zeta^{\prime 4}})p^{2},$
$\tilde{H}_{\text{SXPa}}$ $=\frac{\tilde{\omega}_{c}}{4}(1-\frac{16}{\mu^{4}})x^{2}-\frac{\tilde{\omega}_{c}}{2}(1+\frac{4}{\mu^{2}})-\tilde{\omega}_{0}(\frac{\mu^{2}}{16}+\frac{1}{\mu^{2}}),$
$\tilde{H}_{\text{SXPb}}$ $=\frac{\tilde{\omega}_{c}}{4}(1-\frac{16}{\lambda^{4}})x^{2}-\frac{\tilde{\omega}_{c}}{2}(1-\frac{4}{\lambda^{2}})-\tilde{\omega}_{0}(\frac{\lambda^{2}}{16}+\frac{1}{\lambda^{2}}).$
@@
$\begin{array}[]{r@{}l}\omega_{\text{SX}}&{}=\tilde{\omega}_{c}\sqrt{(1-\frac{1}{\zeta^{4}})(1-\frac{\zeta^{\prime 2}}{\zeta^{2}})},\\ \omega_{\text{SP}}&{}=\tilde{\omega}_{c}\sqrt{(1-\frac{1}{\zeta^{\prime 4}})(1-\frac{\zeta^{2}}{\zeta^{\prime 2}})},\\ \omega_{\text{SXPa}}&{}=0,\\ \omega_{\text{SXPb}}&{}=0.\end{array}$ (32)
We find that the excitation energy closes both in SXPa and SXPb phases. This result denotes the existence of the Goldstone modes Hwang2016JC ; Goldstone . Furthermore, $\omega_{\text{{SP}}}$ has the same form as $\omega_{\text{{SX}}}$ if making the replacement $\zeta\leftrightarrow$$\zeta^{\prime}$ in $\omega_{\text{{SX}}}$.
[…]
V Phase diagram
Figure 5: (Color online) $\langle x\rangle$ in panel (a) and $\langle p\rangle$ in panel (b) versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ for $\eta=10^{2}$.
To study the phase diagram, we introduce the average values $\langle x\rangle$ and $\langle p\rangle$ as the order parameters. The average values $\langle x\rangle$ and $\langle p\rangle$, and the corresponding variances $\Delta x^{2}$ and $\Delta p^{2}$ for p phase are obtained from the ground state of Hamiltonian $\tilde{H}_{p}$ (p=N, SX, SP, SXPa, SXPb). We show the results in table 1. The expressions of $s$, $v$ and $w$ in table 1 are defined as
[…]
$v(x_{1},x_{2})$ $=\frac{1}{2}\sqrt{\frac{1-x_{2}^{2}}{1-x_{1}^{2}}},$ (33b)
$w(x_{1},x_{2})$ $=\frac{\sqrt{x_{1}}}{1+x_{1}}\frac{x_{2}^{2}}{\sqrt{x_{2}^{4}-1}}.$ (33c)
We find that, in the limit $\tilde{\eta}\rightarrow\infty$, $\langle x\rangle\rightarrow\infty$ and $\langle p\rangle=0$ in SX phase while $\langle x\rangle=0$ and $\langle p\rangle\rightarrow\infty$ in the SP phase. That is why they are defined as SX phase and SP phase, respectively. We also note that, the variances $\Delta x^{2}$ and $\Delta p^{2}$ are finite for N, SX and SP phases, but $\Delta x^{2}=0$ and $\Delta p^{2}=\infty$ for both SXPa and SXPb phases.
Table 1: $\langle x\rangle$, $\langle p\rangle$, $\langle x^{2}\rangle$ and $\langle p^{2}\rangle$ for all phases.
@@ V Phase diagram
To see clearly the dependence of $\langle x\rangle$ and $\langle p\rangle$ on the coupling strengths $g_{r}$ and $g_{cr}$, we plot $\langle x\rangle$ and $\langle p\rangle$ versus the coupling strengths $g_{r}$ and $g_{cr}$ in Fig. 5. In the region $g_{r}<\left|\tilde{g}_{C}\right|$, both $\langle x\rangle$ and $\langle p\rangle$ are 0, which correspond to the normal phase. Outside the normal phase region, i.e., $g_{r}>\left|\tilde{g}_{C}\right|$, when $\epsilon>0$, the ground state is in SX phase with $\langle x\rangle>0$; when $\epsilon<0$, the ground state is in SP phase with $\langle p\rangle>0$. One can distinguish the phases by $\langle x\rangle$ and $\langle p\rangle$.
Figure 6: (a) (Color online) (a-b) The first derivatives $\partial E_{G}/\partial\lambda$ and $\partial E_{G}/\partial\mu$ versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$. (b-c) The second derivatives $\partial^{2}E_{G}/\partial\lambda^{2}$ and $\partial^{2}E_{G}/\partial\mu^{2}$ versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$. Other parameters are $\omega_{0}=1,\eta=10^{2}$.

### v2

@@ IV Excitation spectra
To figure out the ground-state property of the system in the entire coupling-parameter space, we calculate the excitation energy in the limit $\tilde{\eta}\rightarrow\infty$ in this section.
@@ IV.1 Normal phase
Then, in the limit $\tilde{\eta}\rightarrow\infty$, we make a unitary transformation $U_{\text{N}}^{\dagger}{H}_{\text{AR}}U_{\text{N}}$ on Hamiltonian (17) with
@@
The transformed Hamiltonian becomes free of coupling between qubit subspaces $\mathcal{H}_{g}$ and $\mathcal{H}_{e}$, which are spanned by $\left|g\right\rangle$ and $\left|e\right\rangle$, respectively. After a projection into $\mathcal{H}_{g}$, i.e., $\tilde{H}_{\text{N}}=\langle g|U_{\text{N}}^{\dagger}H_{\text{AR}}U_{\text{N}}|g\rangle$, we obtain an effective low-energy Hamiltonian
@@
upto a constant term, where
$\zeta=\frac{\lambda(1+\epsilon)}{2},\hskip 14.22636pt\zeta^{\prime}=\frac{\lambda(1-\epsilon)}{2},$ (21)
with the dimensionless rotating-coupling strength $\lambda\equiv{g_{r}/g_{C}}$. We note that there is no interaction term in Eq. (20). From Eq. (20), The excitation energy $\omega_{\text{N}}$ can be directly obtained as
@@
$\tilde{H}_{\text{N}}$ implies divergent values of $\langle x\rangle$ ($\langle p\rangle$) due to the disappearance of $x^{2}$ ($p^{2}$) term when $\epsilon>0$ ($\epsilon<0$) in the limit $\tilde{\eta}\rightarrow\infty$. Due to this fact, the normal phase is restricted in the region
@@
In the normal phase, the value of $\langle x\rangle\ (\langle p\rangle)$ is zero. $\tilde{g}_{C}$ determines the phase transition boundary for the anisotropic QRM. Note that, $g_{C}$ is the critical coupling strength for the standard QRM. When $g_{r}<\left|\tilde{g}_{C}\right|$, the ground state is in the normal phase with the cavity field in a vacuum and qubit in $\left|g\right\rangle$. While outside the normal phase, namely, $g_{r}>\left|\tilde{g}_{C}\right|$, the system is in the superradiant phase with both the qubit and the cavity field being excited.
In Fig. 3, we plot $\tilde{g}_{C}/\omega_{0}$ versus $\xi$ for $\nu/\omega_{0}=0.402$, $0.68$, and $0.8$. For $\nu/\omega_{0}=0.402$, $\tilde{g}_{C}/\omega_{0}$ oscillates between 0 and 0.031. The zero points of $\tilde{g}_{C}/\omega_{0}$ are located at $\xi=5.136$, $8.417$, and $11.62$, corresponding to the zeros of $J_{n_{0}}(\xi)$ ($n_{0}=-2$). For $(1+1/\eta)/(\left|n\right|+3/2)\leqslant\nu/\omega_{0}<(1-1/\eta)/(\left|n\right|+1/2)$ with $n\in\left\{-N\right\}$ and $\nu/\omega_{0}\geq 2(1+1/\eta)$, $\tilde{g}_{C}/\omega_{0}$ becomes independent of $\xi$ due to $m_{0}=n_{0}$. For example, when $\eta=10^{2}$ and $n=0$, the value of $\nu/\omega_{0}$ is between 0.673 and 1.98, such as $\nu/\omega_{0}=0.68$ and $0.8$ shown in Fig. 3. When $m_{0}=n_{0}$, the QPT boundary $\tilde{g}_{C}$ of the anisotropic QRM returns back to the critical point ${g}_{C}$ of the QRM. Thus, the QPT boundary can be tuned via $\nu$ and $\xi$.
@@ IV.2 Superradiant phase
Now, we study the excitation energy in the superradiant phase, namely, $g_{r}>\left|\tilde{g}_{C}\right|$. In the superradiant phase, there are four different phases. When $\epsilon>0$, the superradiant phase can be defined as superradiant $X$ (SX) phase since the $\langle x\rangle$ becomes divergent. When $\epsilon<0$, the system is in superradiant $P$ (SP) phase since the $\langle p\rangle$ is divergent. For $g_{cr}>2\left|g_{C}\right|$ with $g_{r}=0$, the ground state of system is in superradiant XP $a$ (SXPa) phase while for $g_{r}>2\left|g_{C}\right|$ with $g_{cr}=0$, the system is in superradiant XP $b$ (SXPb) phase. When $g_{cr}=0$, only rotating terms exist and the Hamiltonian reduces to a Jaynes-Cummings model. In the following, we investigate the excitation energy in the limit $\tilde{\eta}\rightarrow\infty$ for SX, SP, SXPa, and SXPb phase, respectively.
To effectively describe the superradiant phase with macroscopic excitation, we displace the cavity field $a$ by the displacing operator $D_{k}(\alpha)$ ($k$=SX, SP, SXPa, SXPb) for the $k$ phase. After the displacement, the Hamiltonian $H_{\text{AR}}$ becomes
@@
Here, the detailed form of $D_{k}(\alpha)$ is given in the Appendix B. To diagonalize the qubit part in Hamiltonian $\tilde{H}_{AR}$, we introduce a new set of Pauli operators $\tau_{x,y,z}^{k}$ ($k$=SX, SP, SXPa, SXPb) for each k phase, respectively. To be concise, the form of $\tau_{x,y,z}^{k}$ for the $k$ phase is also given in the Appendix B. In terms of the new Pauli operators $\tau_{x,y,z}^{k}$ , the displaced Hamiltonian $\tilde{H}_{\text{AR}}$ becomes
@@
$\tilde{H}_{\text{AR}}(\alpha_{\text{SXPa}})=$ $\frac{\Omega_{\text{SXPa}}}{2}\tau_{z}^{{}_{\text{SXPa}}}+\chi_{3}\left(\frac{4}{\mu}\tau_{x}^{{}_{\text{SXPa}}}x+\mu\tau_{y}^{{}_{\text{SXPa}}}p\right)$
@@
$\tilde{H}_{\text{AR}}(\alpha_{\text{SXPb}})=$ $\frac{\Omega_{\text{SXPb}}}{2}\tau_{z}^{{}_{\text{SXPb}}}+\chi_{3}\left(\frac{4}{\lambda}\tau_{x}^{{}_{\text{SXPb}}}x-\lambda\tau_{y}^{{}_{\text{SXPb}}}p\right)$
@@
and the rescaled qubit frequency $\Omega_{k}$ for the $k$ phase is
@@
The displacement parameter $\alpha$ is determined by the vanishing of the non-diagonal terms during deriving Eq. (26), which is obtained as
$\alpha_{\text{SX}}$ =±$=\pm\sqrt{\frac{\tilde{\eta}}{2}\left(\zeta^{2}-\frac{1}{\zeta^{2}}\right)},$ (29)
$\alpha_{\text{SP}}$ =±$=\pm\sqrt{\frac{\tilde{\eta}}{2}\left(\zeta^{\prime 2}-\frac{1}{\zeta^{\prime 2}}\right)},$
$\alpha_{\text{SXPa}}$ =±$=\pm\sqrt{\tilde{\eta}\left(\frac{\mu^{2}}{16}-\frac{1}{\mu^{2}}\right)},$
$\alpha_{\text{SXPb}}$ =±$=\pm\sqrt{\tilde{\eta}\left(\frac{\lambda^{2}}{16}-\frac{1}{\lambda^{2}}\right)},$
for the $k$ ($k$=SX, SP, SXPa, SXPb) phase. To distinguish $\alpha$ in each $k$ phase, here we add a subscript $k$ in $\alpha$, i.e., we label $\alpha$ as $\alpha_{k}$. Here, the dimensionless CR coupling strength $\mu$ is defined as $\mu=g_{cr}/g_{C}$. The parameter $\chi_{i}$ ($i$=1,2,3) in Eq. (26) is defined as
@@
Thus, we can make a similar unitary transformation $U_{k}$ as $U_{\text{N}}$ in the normal phase on $\tilde{H}_{\text{AR}}(\alpha_{k})$ to eliminate the non-diagonal terms, namely,
@@
($k$=SX, SP, SXPa, SXPb). Here, the detailed form of $U_{k}$ is given in the Appendix B. After performing a similar procedure used to derive $\tilde{H}_{\text{N}}$, we obtain the effective Hamiltonian $\tilde{H}_{k}$ in the $k$ phase as follows:
$\tilde{H}_{\text{SX}}$ $=\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{1}{\zeta^{4}}\right)x^{2}+\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{\zeta^{\prime 2}}{\zeta^{2}}\right)p^{2},$ (32)
$\tilde{H}_{\text{SP}}$ $=\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{\zeta^{2}}{\zeta^{\prime 2}}\right)x^{2}+\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{1}{\zeta^{\prime 4}}\right)p^{2},$
$\tilde{H}_{\text{SXPa}}$ $=\frac{\tilde{\omega}_{c}}{4}\left(1-\frac{16}{\mu^{4}}\right)x^{2}-\frac{\tilde{\omega}_{c}}{2}\left(1+\frac{4}{\mu^{2}}\right)-\tilde{\omega}_{0}\left(\frac{\mu^{2}}{16}+\frac{1}{\mu^{2}}\right),$
$\tilde{H}_{\text{SXPb}}$ $=\frac{\tilde{\omega}_{c}}{4}\left(1-\frac{16}{\lambda^{4}}\right)x^{2}-\frac{\tilde{\omega}_{c}}{2}\left(1-\frac{4}{\lambda^{2}}\right)-\tilde{\omega}_{0}\left(\frac{\lambda^{2}}{16}+\frac{1}{\lambda^{2}}\right).$
@@
$\begin{array}[]{r@{}l}\omega_{\text{SX}}&{}=\tilde{\omega}_{c}\sqrt{\left(1-\frac{1}{\zeta^{4}}\right)\left(1-\frac{\zeta^{\prime 2}}{\zeta^{2}}\right)},\\ \omega_{\text{SP}}&{}=\tilde{\omega}_{c}\sqrt{\left(1-\frac{1}{\zeta^{\prime 4}}\right)\left(1-\frac{\zeta^{2}}{\zeta^{\prime 2}}\right)},\\ \omega_{\text{SXPa}}&{}=0,\\ \omega_{\text{SXPb}}&{}=0.\end{array}$ (33)
Table 1: $\langle x\rangle$, $\langle p\rangle$, $\langle x^{2}\rangle$, and $\langle p^{2}\rangle$ for all phases.
@@
We find that the excitation energy closes both in the SXPa and SXPb phases. This result denotes the existence of the Goldstone modes Hwang2016JC ; Goldstone . Furthermore, $\omega_{\text{{SP}}}$ has the same form as $\omega_{\text{{SX}}}$ if making the replacement $\zeta\leftrightarrow$$\zeta^{\prime}$ in $\omega_{\text{{SX}}}$.
Figure 5: $\langle x\rangle$ in panel (a) and $\langle p\rangle$ in panel (b) versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ for $\eta=10^{2}$.
[…]
Figure 6: (a), (b) The first derivatives $\partial E_{G}/\partial\lambda$ and $\partial E_{G}/\partial\mu$ versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$. (c), (d) The second derivatives $\partial^{2}E_{G}/\partial\lambda^{2}$ and $\partial^{2}E_{G}/\partial\mu^{2}$ versus the coupling strengths $g_{r}/g_{C}$ and $g_{cr}/g_{C}$. Other parameters are $\omega_{0}=1,\eta=10^{2}$.
V Phase diagram
To study the phase diagram, we introduce the average values $\langle x\rangle$ and $\langle p\rangle$ as the order parameters. The average values $\langle x\rangle$ and $\langle p\rangle$, and the corresponding variances $\Delta x^{2}$ and $\Delta p^{2}$ for $p$ phase are obtained from the ground state of the Hamiltonian $\tilde{H}_{p}$ ($p$=N, SX, SP, SXPa, SXPb). We show the results in Table 1. The expressions of $s$, $v$, and $w$ in Table 1 are defined as
Figure 7: (Color online) The excitation energy $\omega$ versus the dimensionless modulation amplitude $\xi$. Parameters are $\eta=10^{2},\nu/\omega_{0}=0.68$, and $g/\omega_{0}=0.06$.
[…]
$v(x_{1},x_{2})$ $=\frac{1}{2}\sqrt{\frac{1-x_{2}^{2}}{1-x_{1}^{2}}},$ (34b)
$w(x_{1},x_{2})$ $=\frac{\sqrt{x_{1}}}{1+x_{1}}\frac{x_{2}^{2}}{\sqrt{x_{2}^{4}-1}}.$ (34c)
It is found that, in the limit $\tilde{\eta}\rightarrow\infty$, we have $\langle x\rangle\rightarrow\infty$ and $\langle p\rangle=0$ in the SX phase while $\langle x\rangle=0$ and $\langle p\rangle\rightarrow\infty$ in the the SP phase. That is why they are defined as the SX phase and SP phase, respectively. We also note that the variances $\Delta x^{2}$ and $\Delta p^{2}$ are finite for N, SX, and SP phases, but $\Delta x^{2}=0$ and $\Delta p^{2}=\infty$ for both the SXPa and SXPb phases.
To see clearly the dependence of $\langle x\rangle$ and $\langle p\rangle$ on the coupling strengths $g_{r}$ and $g_{cr}$, we plot $\langle x\rangle$ and $\langle p\rangle$ versus the coupling strengths $g_{r}$ and $g_{cr}$ in Fig. 5. In the region $g_{r}<\left|\tilde{g}_{C}\right|$, both $\langle x\rangle$ and $\langle p\rangle$ are 0, which corresponds to the normal phase, as labeled by $N$. Outside the normal phase region, i.e., $g_{r}>\left|\tilde{g}_{C}\right|$, when $\epsilon>0$, the ground state is in the SX phase with $\langle x\rangle>0$; when $\epsilon<0$, the ground state is in the SP phase with $\langle p\rangle>0$. One can distinguish the phases by $\langle x\rangle$ and $\langle p\rangle$.